# PointNSP: Autoregressive 3D Point Cloud Generation with Next-Scale Level-of-Detail Prediction

### v2

@@ Abstract
Autoregressive point cloud generation has long lagged behind diffusion-based approaches in quality. The performance gap stems from the fact that autoregressive models impose an artificial ordering on inherently unordered point sets, forcing shape generation to proceed as a sequence of local predictions. This sequential bias emphasizes short-range continuity but undermines the model’s capacity to capture long-range dependencies, hindering its ability to enforce global structural properties such as symmetry, consistent topology, and large-scale geometric regularities. Inspired by the level-of-detail (LOD) principle in shape modeling, we propose PointNSP, a coarse-to-fine generative framework that preserves global shape structure at low resolutions and progressively refines fine-grained geometry at higher scales through a next-scale prediction paradigm. This multi-scale factorization aligns the autoregressive objective with the permutation-invariant nature of point sets, enabling rich intra-scale interactions while avoiding brittle fixed orderings. Experiments on ShapeNet show that PointNSP establishes state-of-the-art (SOTA) generation quality for the first time within the autoregressive paradigm. In addition, it surpasses strong diffusion-based baselines in parameter, training, and inference efficiency. Finally, in dense generation with 8,192 points, PointNSP ’s advantages become even more pronounced, underscoring its scalability potential. [^1] Project Homepage: https://pointnsp.pages.dev
@@ 1 Introduction
Figure 1: A comprehensive performance comparison between PointNSP and recent strong baseline methods across six key metrics.
Point clouds are a fundamental representation of 3D object shapes, describing each object as a collection of points in Euclidean space. They arise naturally from sensors such as LiDAR and laser scanners, offering a compact yet expressive encoding of fine-grained geometric details. Developing powerful generative models for point clouds is key to uncovering the underlying distribution of 3D shapes, with broad applications in shape synthesis, reconstruction, computer-aided design, and perception for robotics and autonomous systems. However, high-fidelity point cloud generation remains challenging due to the irregular and unordered nature of point sets [PointNet, PointNet++, DeepSet]. Unlike images or sequences, point clouds have no inherent ordering—permuting the points leaves the shape unchanged—making naive order-dependent methods poorly suited for their generation and analysis.
In recent years, diffusion-based methods [PVD, LION, Tiger] have become the dominant paradigm for 3D point cloud generation, delivering strong results. However, their Markovian assumption overlooks global context, often leading to incoherent shapes. Moreover, diffusion models are computationally costly, as generating high-quality samples typically requires hundreds to thousands of denoising steps—a burden that grows prohibitive when producing dense point clouds. By contrast, autoregressive (AR) models condition on the full history, enabling explicit modeling of long-range dependencies while generally offering faster sampling. Unlike diffusion models, existing AR methods require flattening inherently unordered point clouds into sequences. Point Transformers [PointTransformer, PointTransformerV2, PointTransformerV3] design specialized architectures for unordered point sets and explore diverse point cloud serialization strategies to improve speed, at the expense of relaxing permutation invariance. PointGrow [PointGrow] enforces a sequential order by sorting points along the $z$-axis. ShapeFormer [shapeformer] voxelizes point clouds and flattens codebook embeddings into sequences using a row-major order. PointVQVAE [pointvqvae] projects patches onto a sphere and arranges them in a spiral sequence to establish a canonical mapping. AutoSDF [AutoSDF] treats point clouds as randomly permuted sequences of latent variables, while PointGPT [PointGPT] leverages Morton code ordering to impose structure on unordered data. Although these approaches yield promising results, they still lag behind state-of-the-art diffusion-based methods [DPM, PVD, LION, Tiger] in generation quality. This is largely due to the restrictive unidirectional dependencies imposed by fixed sequential orders, which collapse global shape generation into local predictions and inherently violate the fundamental permutation-invariance property. This naturally raises the question: can we achieve permutation-invariant autoregressive modeling for 3D point cloud generation?
@@
In this work, we introduce PointNSP, a novel autoregressive framework for 3D point cloud generation that preserves global permutation invariance—a key property ensuring that shapes remain independent of point ordering. Inspired by visual autoregressive modeling in 2D image synthesis [VAR], PointNSP follows a coarse-to-fine strategy, progressively refining point clouds from global structures to fine-grained details via next-scale prediction. Unlike prior approaches that predict one point at a time (next-point prediction), PointNSP captures multiple levels of detail (LoD) [LoD] at each step, enabling more effective modeling of both global geometry and local structure. This design offers two key advantages. First, it avoids collapsing 3D structures by eliminating the need to flatten point clouds into 1D sequences: each step corresponds to a full 3D shape at a given LoD, ensuring structural coherence and permutation invariance. Second, compared to diffusion-based methods, PointNSP establishes a more structured and efficient generation trajectory, avoiding iterative noise injection and denoising in 3D space. Together, these advances allow PointNSP to achieve high generation quality while maintaining strong modeling efficiency. The comparisons across different paradigms are illustrated in Figure 2.
@@
We conduct extensive experiments on the ShapeNet benchmark to validate the effectiveness of PointNSP across diverse settings. In the standard single-class scenario, PointNSP achieves state-of-the-art (SOTA) generation quality, yielding the lowest average Chamfer Distance and Earth Mover’s Distance—setting a new benchmark for autoregressive modeling. Beyond quality, PointNSP also demonstrates substantially higher parameter efficiency, training efficiency, and sampling speed compared to strong diffusion-based baselines. In the more challenging multi-class setting, PointNSP maintains SOTA performance, evidencing superior cross-category generalization. Moreover, PointNSP significantly outperforms existing approaches on downstream tasks such as partial point cloud completion and upsampling, further highlighting the robustness and versatility of its design. Comparative results across these metrics are presented in Figure 1. When evaluated on denser point clouds with 8192 points, the advantages of PointNSP become even more pronounced, particularly in the aforementioned efficiency metrics, underscoring its scalability potential.
@@ 2 Related Works
Deep generative models have made significant strides in 3D point cloud generation. For instance, PointFlow [PointFlow] captures the latent distribution of point clouds using continuous normalizing flows [normalizing-flow, ffjord, neuralode]. Building on this foundation, a series of approaches—including DPM [DPM], ShapeGF [ShapeGF], PVD [PVD], LION [LION], TIGER [Tiger], PDT [wang2025pdt]—leverage denoising diffusion probabilistic models [ddpm, score-sde] to synthesize 3D point clouds through the gradual denoising of input data or latent representations in continuous space. While efforts to improve the sampling efficiency of diffusion models—such as straight flows [flow-matching, rectified-flow, PSF] and ODE-based solvers [DDIM]—have achieved acceleration, these methods often introduce trade-offs that compromise generation quality. In contrast, autoregressive models for 3D point cloud generation [PointGrow, pointvqvae, PointGPT] have received relatively less attention, historically lagging behind diffusion-based techniques in terms of fidelity. In this work, we reformulate point cloud generation as an iterative upsampling process, establishing a connection with sparse point cloud upsampling approaches [PU-net, Grad-PU, PUDM]. A more comprehensive review of related point cloud upsampling methods is provided in the Appendix F.
@@ 3.1 Preliminaries: Autoregressive 3D Point Cloud Generation
Training Eq. 1 requires a sequential classification objective. Each point $\mathbf{x}_{i}$ is first converted into a discrete integer token $q_{i}$ via VQ-VAE quantization [VQVAE], and the resulting tokens are flattened into a sequence $(q_{1},\dots,q_{N})$ according to a predefined generation order. Autoregressive modeling is then applied to this discrete sequence following the paradigm in Eq. 1. However, this approach struggles to preserve permutation invariance, since the probability distribution depends on the chosen token ordering and is not invariant to different permutations of the points. In this work, instead of predicting the next point, we propose a novel autoregressive framework that predicts the next-scale level-of-detail (LoD) of the point cloud $\mathbf{X}$, while simultaneously preserving the permutation invariance property:
@@
In this work, we set $s_{1}=1$ (a single starting point) and $s_{K}=N$ (total number of points). The generation process can be interpreted as an autoregressive upsampling procedure with sequential upsampling rates $r_{1},r_{2},\dots,r_{K-1}$, satisfying the relationship $r_{K-1}\times\dots\times r_{1}\times s_{1}=s_{K}=N$. Learning Eq. 3 requires a two-stage training process. The first stage involves training multi-scale VQ-VAE tokenizers to progressively reconstruct different LoDs, while the second stage trains the autoregressive transformer over the resulting multi-scale token sequence.
@@
6:  $q_{k}=\mathcal{Q}(f_{k}),f_{k}=\operatorname{Query}(f,\mathbf{X}_{k})$
@@
12:Return: token sequence $Q$.
@@
1:Input: token sequence $Q$.
@@
3.2 Multi-Scale LoD Representation
Sampling Multi-Scale LoD Sequence.
To obtain $\mathbf{X}_{1},\dots,\mathbf{X}_{K}$}, we employ the Farthest Point Sampling ($\operatorname{FPS}$[FPS] algorithm to downsample the original point cloud $\mathbf{X}$. $\operatorname{FPS}$ offers two key advantages. First, it is permutation-invariant: point selection depends solely on spatial locations (geometric distances) rather than the order of points in the input array. The stochasticity of the downsampled point cloud arises only from the choice of the initial point, not from the input ordering. Second, $\operatorname{FPS}$ operates directly on the original point cloud, enabling straightforward querying of specific point embeddings via stored indices. Importantly, we do not fix the initial starting point of $\operatorname{FPS}$, so the sampled points in $\mathbf{X}_{k}$ vary during training. This inherent stochasticity enhances model generalization, effectively serving as a form of data augmentation at each scale. Note that $\mathbf{X}_{1},\dots,\mathbf{X}_{K}$} are all obtained directly from the original point cloud $\mathbf{X}$, i.e., $\mathbf{X}_{k}=\operatorname{FPS}(\mathbf{X})$ for each scale $k$, rather than using an iterative downsampling strategy such as $\mathbf{X}_{k}=\operatorname{FPS}^{k}(\mathbf{X})$ adopted in 2D image VAR [VAR], where $\operatorname{FPS}^{k}$ denotes a nested operation with $k$ loops. Directly sampling points from $\mathbf{X}$ in this way increases the shape coverage at each scale of the LoD.
Multi-Scale Feature Extraction.
For the encoder to obtain point features $f\in\mathbb{R}^{N\times 3}$ from the point cloud $\mathbf{X}$, any permutation-equivariant network $\text{NN}(\cdot)$ is applicable: the per-point features reorder consistently with any permutation of the input points, i.e., $\pi(\text{NN}(\mathbf{x}_{1},...,\mathbf{x}_{N}))=\text{NN}(\pi(\mathbf{x}_{1},...,\mathbf{x}_{N}))$ for any permutation $\pi$. Therefore, permutation-equivariant architectures such as PointNet [PointNet], PointNet++ [PointNet++], PointNeXt [PointNext] and PVCNN [PV-CNN] are all applicable. To construct the multi-scale representations, we extract latent features $f_{1}\in\mathbb{R}^{s_{1}\times d},\dots,f_{K}\in\mathbb{R}^{s_{K}\times d}$ in a residual querying manner: each $\mathbf{X}_{k}$ retrieves features from the portion of $f$ not already captured by earlier scales, using the stored indices from downsampling. This ensures that later scales focus on complementary information rather than duplicating what has already been encoded. The resulting features are then used for multi-scale VQVAE codebook training.
Figure 3: Illustration of training a multi-scale VQ-VAE for 3D point cloud reconstruction across scales $s_{1}$ to $s_{4}$, resulting in a multi-scale token sequence $Q=(q_{1},\dots,q_{4})$.
Multi-Scale RVQ Tokenizer.
We use residual vector quantization (RVQ) [VAR, RQtransformer] to approximate $f$ progressively. Each residual feature $f_{k}$ is quantized into tokens $q_{k}=\{q^{1}_{k},q^{2}_{k},\dots,q^{s_{k}}_{k}\}=\mathcal{Q}(f_{k})\in[V]^{s_{k}}$, where each $q^{i}_{k}$ indexes the nearest codebook embedding $\mathbf{z}_{v}\in\mathbb{R}^{d}$: $q^{i}_{k}=\operatorname*{arg\,min}_{v\in[v]}\lVert\mathbf{z}_{v}-f^{i}_{k}\rVert_{2}$. This produces scale token sets $q_{1},\dots,q_{K}$} and corresponding embeddings $\mathbf{z}_{1}\in\mathbb{R}^{s_{1}\times d},\dots,\mathbf{z}_{K}\in\mathbb{R}^{s_{K}\times d}$}, which together provide a residual approximation of the original feature $f$ as follows:
$f_{k}=f_{k-1}-\phi_{k}(\operatorname{upsampling}(\mathbf{z}_{k},s_{K})),\ f_{0}=f$ (4)
Here $\phi_{k}(\cdot):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times d}$ refines the latent embedding. The $\text{upsampling}(\cdot,s_{K})$ module consists of a series of reshaping operations to upsample the latent $\mathbf{z}_{k}\in\mathbb{R}^{s_{k}\times d}$ to the highest (target) resolution $s_{K}\times d$. Compared to the vanilla VQ [VQVAE], the above RVQ mechanism allows finer approximation of the original $f$ with more efficient codebook utilization. We could compute the partial sum $\hat{f}$ and reconstruct the final predicted 3D shape $\mathbf{\hat{X}}$ as follows:
$\mathbf{\hat{X}}=D(\hat{f}),\ \hat{f}=\sum_{k=1}^{K}\phi_{k}(\operatorname{upsampling}(\mathbf{z}_{k},s_{K})),\ \mathbf{z}_{k}=\operatorname{lookup}(q_{k}),$ (5)
where $D(\cdot):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times 3}$ is a simple MLP decoder. Eq. 5 shows that $\mathbf{\hat{X}}$ is obtained by aggregating information across all scales of the level-of-detail (LoD). The pseudo-algorithms and process are illustrated in Algorithm 1, Algorithm 2, and Figure 3.
Upsampling & Reconstruction.
The upsampling operation in both Eq. 4 and Eq. 5 follows a PU-Net [PU-net] like operation, which consists of the duplicating and reshaping:
$\mathbf{z}_{k}(s_{k}\times d)\xrightarrow{\text{duplicate}}\mathbf{z}_{k}(s_{k}\times r\times d)\xrightarrow{\text{reshape}}\mathbf{z}_{k}((s_{k}*r)\times d)=\mathbf{z}_{K}(s_{K}\times d),$ (6)
where the upsampling rate $r=\frac{s_{K}}{s_{k}}$ can be pre-computed. This upsampling operation can densify points by arbitrary factors while preserving the fundamental permutation-equivariance property. The reconstruction loss function to be minimized is defined as follows:

### v3

@@ Abstract
Autoregressive point cloud generation has long lagged behind diffusion-based approaches in quality. The performance gap stems from the fact that autoregressive models impose an artificial ordering on inherently unordered point sets, forcing shape generation to proceed as a sequence of local predictions. This sequential bias reinforces short-range continuity but limits the model’s ability to capture long-range dependencies, thereby weakening its capacity to enforce global structural properties such as symmetry, geometric consistency, and large-scale spatial regularities. Inspired by the level-of-detail (LOD) principle in shape modeling, we propose PointNSP, a coarse-to-fine generative framework that preserves global shape structure at low resolutions and progressively refines fine-grained geometry at higher scales through a next-scale prediction paradigm. This multi-scale factorization aligns the autoregressive objective with the permutation-invariant nature of point sets, enabling rich intra-scale interactions while avoiding brittle fixed orderings. Strictly following the baseline experimental setups, empirical results on ShapeNet benchmark demonstrate that PointNSP achieves state-of-the-art (SOTA) generation quality for the first time within the autoregressive paradigm. Moreover, it surpasses strong diffusion-based baselines in parameter, training, and inference efficiency. Finally, under dense generation with $8,192$ points, PointNSP’s advantages become even more pronounced, highlighting its strong scalability potential. [^1] Project Homepage: https://pointnsp.pages.dev
@@ 1 Introduction
Figure 1: PointNSP achieves SoTA performance compared to recent strong baseline methods across six key evaluation metrics.
Point clouds are a fundamental representation of 3D object shapes, describing each object as a collection of points in Euclidean space. They arise naturally from sensors such as LiDAR and laser scanners, offering a compact yet expressive encoding of fine-grained geometric details. Developing powerful generative models for point clouds is key to uncovering the underlying distribution of 3D shapes, with broad applications in shape synthesis, reconstruction, computer-aided design, and perception for robotics and autonomous systems. However, high-fidelity point cloud generation remains challenging due to the irregular and unordered nature of point sets [PointNet, PointNet++, DeepSet]. Unlike images or sequences, point clouds lack an inherent ordering—permutations do not alter the underlying shape—rendering naive order-dependent modeling strategies fundamentally misaligned with their structure.
@@
In recent years, diffusion-based methods [PVD, LION, Tiger] have become the dominant paradigm for 3D point cloud generation, achieving strong empirical performance. However, their Markovian formulation prevents them from leveraging full historical context, often resulting in incoherent shapes. They are also computationally expensive: producing high-quality samples typically requires hundreds to thousands of denoising steps, a burden that becomes prohibitive for dense point clouds. Although advanced samplers [DDIM, DPM-solver] can accelerate sampling, such speedups commonly degrade generation quality. By contrast, autoregressive (AR) models condition on the entire generation history, which helps enforce local smoothness and generally enables faster sampling. Yet, existing AR approaches must impose an artificial ordering on the inherently unordered point set, typically by flattening it into a sequence, which adheres to a next-point prediction paradigm. Point Transformers [PointTransformer, PointTransformerV2, PointTransformerV3] design specialized architectures for unordered point sets and explore diverse point cloud serialization strategies to improve speed, at the expense of relaxing permutation invariance. PointGrow [PointGrow] enforces a sequential order by sorting points along the $z$-axis. ShapeFormer [shapeformer] voxelizes point clouds and flattens codebook embeddings into sequences using a row-major order. PointVQVAE [pointvqvae] projects patches onto a sphere and arranges them in a spiral sequence to establish a canonical mapping. AutoSDF [AutoSDF] treats point clouds as randomly permuted sequences of latent variables, while PointGPT [PointGPT] leverages Morton code ordering to impose structure on unordered data. Although these approaches yield promising results, they still lag behind strong diffusion-based methods [DPM, PVD, LION, Tiger] in generation quality. This is largely due to the restrictive unidirectional dependencies imposed by fixed sequential orders, which collapse global shape generation into local predictions and inherently violate the fundamental permutation-invariance property. This naturally raises the question: can we achieve permutation-invariant autoregressive modeling for 3D point cloud generation?
@@
In this work, we introduce PointNSP, a novel autoregressive framework for 3D point cloud generation that preserves global permutation invariance—a key property ensuring that shapes remain independent of point ordering. PointNSP follows a coarse-to-fine strategy, progressively refining point clouds from global structures to fine-grained details via next-scale prediction. Unlike prior approaches that predict one point at a time (next-point prediction), PointNSP captures multiple levels of detail (LoD) [LoD] at each step, enabling more effective modeling of both global geometry and local structure. This design offers two key advantages. First, it avoids collapsing 3D structures by eliminating the need to flatten point clouds into 1D sequences: each step corresponds to a full 3D shape at a given LoD, ensuring structural coherence and permutation invariance. Second, compared to diffusion-based methods, PointNSP establishes a more structured and efficient generation trajectory, avoiding iterative noise injection and denoising in 3D space. Together, these advances allow PointNSP to achieve high generation quality while maintaining strong modeling efficiency. The comparisons across different paradigms are illustrated in Figure 2.
@@
We conduct extensive experiments on the ShapeNet benchmark to validate the effectiveness of PointNSP across diverse settings. In the standard single-class scenario, PointNSP achieves state-of-the-art (SoTA) generation quality, yielding the lowest average Chamfer Distance and Earth Mover’s Distance—setting a new benchmark for autoregressive modeling. Beyond quality, PointNSP also demonstrates substantially higher parameter efficiency, training efficiency, and sampling speed compared to strong diffusion-based baselines. In the more challenging many-class ($55$-class) generation setting, PointNSP maintains SoTA performance, evidencing superior cross-category generalization. Moreover, PointNSP significantly outperforms existing approaches on downstream tasks such as partial point cloud completion and upsampling, further highlighting the robustness and versatility of its design. Comparative results across these metrics are presented in Figure 1. When evaluated on denser point clouds with $8192$ points, the advantages of PointNSP become even more pronounced, particularly in the aforementioned efficiency metrics, underscoring its scalability potential.
@@ 2 Related Works
Deep generative models have made significant strides in 3D point cloud generation. For instance, PointFlow [PointFlow] captures the latent distribution of point clouds using continuous normalizing flows [normalizing-flow, ffjord, neuralode]. Building on this foundation, a series of approaches—including DPM [DPM], ShapeGF [ShapeGF], PVD [PVD], LION [LION], TIGER [Tiger], PDT [wang2025pdt]—leverage denoising diffusion probabilistic models [ddpm, score-sde] to synthesize 3D point clouds through the gradual denoising of input data or latent representations in continuous space. While efforts to improve the sampling efficiency of diffusion models—such as straight flows [flow-matching, rectified-flow, PSF] and ODE-based solvers [DDIM]—have achieved acceleration, these methods often introduce trade-offs that compromise generation quality. In contrast, autoregressive models for 3D point cloud generation [PointGrow, pointvqvae, PointGPT] have received relatively less attention, historically lagging behind diffusion-based techniques in terms of fidelity. In this work, we reformulate point cloud generation as an iterative upsampling process, establishing a connection with sparse point cloud upsampling approaches [PU-net, Grad-PU, PUDM]. A comprehensive review of point cloud upsampling works is provided in the Appendix 11.
@@ 3.1 Preliminaries: Autoregressive 3D Point Cloud Generation
Training Eq. 1 requires a sequential classification objective. Each point $\mathbf{x}_{i}$ is first converted into a discrete integer token $q_{i}$ via VQ-VAE quantization [VQVAE], and the resulting tokens are flattened into a sequence $(q_{1},\dots,q_{N})$ according to a predefined generation order. Autoregressive modeling is then applied to this discrete sequence following the paradigm in Eq. 1. However, this approach struggles to preserve permutation invariance, since the probability distribution depends on the chosen token ordering and is not invariant to different permutations of the points.
Figure 3: (a) Illustration of training a multi-scale VQVAE in a residual manner for point cloud representation across scales $s_{1}$ to $s_{3}$, resulting in a multi-scale token sequence $Q=(q_{1},\dots,q_{3})$; (b) Illustration of training a causal transformer with intermediate shape decoding, scale token $\operatorname{upsampling}$ ($s_{1}\rightarrow s_{2}$ and $s_{2}\rightarrow s_{3}$), position-aware soft masks $\mathbf{M}^{P}_{k}$, and block-wise causal masks $\mathbf{M}$.
In this work, instead of predicting the next point, we propose a novel autoregressive framework that predicts the next-scale level-of-detail (LoD) of the point cloud $\mathbf{X}$, while simultaneously preserving the permutation invariance property:
@@
The generation process bears a strong resemblance to an autoregressive upsampling procedure, governed by a sequence of upsampling rates $r_{1},r_{2},\dots,r_{K-1}$ that satisfy the relation $r_{K-1}\times\dots\times r_{1}\times s_{1}=s_{K}$. We set $s_{K}=N,\mathbf{X}_{K}=\mathbf{X}$ to reconstruct the shape at the target full-resolution point count. Under this formulation, Eq. 3 requires learning a hierarchy of representations $\mathbf{X}_{k}$ across multiple LoD information. These representations progressively encode more global and semantically coherent structural information than the next-point prediction in Eq. 1.
3.2 Multi-Scale LoD Representation
Sampling LoD Sequence.
To construct the LoD sequence $(\mathbf{X}_{1},\dots,\mathbf{X}_{K})$ for each sample $\mathbf{X}$, three key considerations must be addressed. First, the sampling operation should preserve the permutation invariance property of point clouds, ensuring that the resulting LoD sequence remains independent of the input point ordering in $\mathbf{X}$. Second, the sampling strategy should aim for comprehensive spatial coverage at each LoD, reflecting the existence of an underlying continuous surface that the 3D shape represents. To this end, we adopt the Farthest Point Sampling (FPS) [FPS] algorithm to construct the LoD sequence. The point selection mechanism in FPS relies exclusively on spatial geometry—specifically, pairwise Euclidean distances—rather than the ordering of points within the input set, thereby preserving permutation invariance. Moreover, the inherent stochasticity introduced by the random initialization of the starting point enables diverse subsets $\mathbf{X}_{k}$ to be sampled at each scale $k$ across different training epochs, thereby enhancing the diversity and representational robustness of the generated LoD sequences.
Multi-Scale Feature Extraction.
Rather than directly learning tokenizers from the sampled point clouds in the 3D coordinate space, we learn their corresponding quantized representations within the latent feature space defined by the LoD sequence. To obtain latent features $\mathbf{f}^{0}\in\mathbb{R}^{N\times d}$ from the point cloud $\mathbf{X}$, any permutation-equivariant network $\text{NN}(\cdot)$ is applicable: the per-point features reorder consistently with any permutation of the input points, i.e., $\pi(\text{NN}(\mathbf{x}_{1},...,\mathbf{x}_{N}))=\text{NN}(\pi(\mathbf{x}_{1},...,\mathbf{x}_{N}))$ for any permutation $\pi$. Therefore, permutation-equivariant architectures such as PointNet [PointNet], PointNet++ [PointNet++], PointNeXt [PointNext] and PVCNN [PV-CNN] are all applicable. To encourage each scale to capture complementary information rather than redundantly encoding features already represented at coarser levels, we extract latent features $\mathbf{f}_{k}\in\mathbb{R}^{s_{k}\times d}$ in a residual fashion:
$\mathbf{f}_{k}=\operatorname{query}(\mathbf{f}^{k-2}-\tilde{\mathbf{f}}_{k-1},\mathbf{X}_{k}),\ \mathbf{f}_{1}=\operatorname{query}(\mathbf{f}^{0},\mathbf{X}_{1}).$ (4)
Here, $\tilde{\mathbf{f}}_{k-1}$ represents the feature contribution from the learned tokenizers at scale k1k-1, while $\operatorname{query}(\cdot)$ retrieves latent features according to the index correspondence between the sampled subset $\mathbf{X}_{k}$ and the original point set $\mathbf{X}$, ensuring consistent alignment across scales. This produces a sequence of LoD latent features $(\mathbf{f}_{1},\dots,\mathbf{f}_{K})$.
Multi-Scale VQVAE Tokenizer.
For each scale $k$, we learn tokenizers for latent feature $\mathbf{f}_{k}$ through quantization $\mathcal{Q}$ into discrete tokens $q_{k}=(q^{1}_{k},q^{2}_{k},\dots,q^{s_{k}}_{k})=\mathcal{Q}(\mathbf{f}_{k})\in[V]$, where $[V]$ denotes a sequence of entries corresponding to indices in a learnable codebook $Z\in\mathbb{R}^{V\times d}$ containing $V$ vectors. Note that this codebook $Z$ is shared across all scales for efficient utilization. Each token $q^{i}_{k}$ indexes the nearest codebook embedding $\mathbf{z}_{v}\in\mathbb{R}^{d}$ to the corresponding latent feature $\mathbf{f}_{k}[i]\in\mathbb{R}^{d}$: $q^{i}_{k}=\operatorname*{arg\,min}_{v\in[V]}\lVert\mathbf{z}_{v}-\mathbf{f}_{k}[i]\rVert_{2}$. This produces scale-wise token sequence $Q=(q_{1},\dots,q_{K})$ along with their corresponding embeddings $(\mathbf{z}_{1}\in\mathbb{R}^{s_{1}\times d},\dots,\mathbf{z}_{K}\in\mathbb{R}^{s_{K}\times d})$, forming a hierarchical discrete representation aligned with the LoD sequence. For each scale $k$, the feature contribution $\tilde{\mathbf{f}}_{k}$ produced by the scale-specific tokenizers $q_{k}$ is then given by:
$\tilde{\mathbf{f}}_{k}=\phi_{k}(\operatorname{upsampling}(\mathbf{z}_{k},s_{K})),\ \mathbf{z}_{k}=\operatorname{lookup}(Z,q_{k}),$ (5)
where $\phi_{k}(\cdot):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times d}$ is a permutation-equivariant network that refines the latent embeddings, and $\operatorname{upsampling}(\cdot,s_{K})$ increases the resolution of the latent $\mathbf{z}_{k}\in\mathbb{R}^{s_{k}\times d}$ to the highest resolution $s_{K}\times d$.
Upsampling & Reconstruction.
[…]
$\mathbf{z}_{k}(s_{k}\times d)\xrightarrow{\operatorname{duplicate}}\mathbf{z}_{k}(s_{k}\times r\times d)\xrightarrow{\operatorname{reshape}}\mathbf{z}_{k}((s_{k}\cdot r)\times d),$ (6)
where $\mathbf{z}_{k}\in\mathbb{R}^{(s_{k}\cdot r)\times d}=\mathbf{z}_{K}\in\mathbb{R}^{s_{K}\times d}$ denotes the upsampled representation at the largest scale $s_{K}$, with upsampling rate $r=\frac{s_{K}}{s_{k}}$ specified in advance. This operation densifies points by arbitrary factors while preserving the permutation-equivariance of the latent features. The reconstruction loss is defined as:
[…]
where $\mathcal{L}_{\text{CD}}$ and $\mathcal{L}_{\text{EMD}}$ denote the Chamfer Distance and Earth Mover’s Distance losses, commonly used to evaluate point cloud similarity from complementary perspectives. The stop-gradient operation $sg[\cdot]$ ensures that the latent features $\mathbf{f}_{k}$ used for reconstruction remain consistent with the quantized latent vectors $\mathbf{z}_{k}$. The above process is described in Algorithm 1 and  2 and Figure 3 (a).
@@ 3.2 Multi-Scale LoD Representation
6:  $q_{k}=\mathcal{Q}(\mathbf{f}_{k}),\mathbf{f}_{k}=\operatorname{query}(\mathbf{f}^{k},\mathbf{X}_{k})$
@@
12:Return: token sequence $Q=(q_{1},\dots,q_{K})$.
@@
1:Input: token sequence $Q=(q_{1},\dots,q_{K})$.
@@
3.3 Autoregressive Transformer for Next-Scale LoD Prediction
The next step is to train an autoregressive transformer on the input multi-scale token sequence $Q=(\operatorname{[start]},q_{1},\dots,q_{K-1})$ to predict $(q_{1},\dots,q_{K})$. Owing to the strong local-geometry inductive bias in 3D structures, a standard causal transformer struggles to capture both intra-scale and inter-scale dependencies efficiently. Therefore, it is necessary to incorporate geometric information into the transformer design, which poses additional challenges due to the unordered nature of point clouds.
Inter-Scale Interaction Modeling.
Inter-scale interactions across input scales $(q_{1},\dots,q_{K-1})$ are critical for the model to capture dependencies between levels of detail and to generate each subsequent LoD conditioned on the preceding ones. We follow a key principle: Tokens at scale $q_{k}$ are permitted to attend only to tokens from preceding scales $q_{1},\dots,q_{k-1}$. Within each scale, however, all tokens $q_{k}=(q^{1}_{k},\dots,q^{s_{k}}_{k})$ are allowed to fully attend to one another, ensuring that the model interprets them as complete shapes at the corresponding resolution. To enforce this constraint, we construct a causal mask $\mathbf{M}\in\mathbb{R}^{(s_{1}+\dots+s_{K})\times(s_{1}+\dots+s_{K})}$ as a block-diagonal matrix, where each diagonal block $\mathbf{M}_{k}$ of size $s_{k}\times s_{k}$ is fully unmasked: $\mathbf{M}=\operatorname{diag}[\mathbf{M}_{1},\mathbf{M}_{2},...,\mathbf{M}_{K}]$. This design allows bidirectional attention within each scale while maintaining a lower-triangular dependency structure across scales, thereby preventing information leakage from future scale tokens. To further distinguish scales, each token is assigned a scale embedding $\mathbf{s}^{i}_{k}\in\mathbb{R}^{d}$, implemented as a one-hot embedding over $K$ scales and shared among tokens within the same scale. Algorithmic details are explained in the Appendix 7.
Intra-Scale Interaction Modeling.
Since the intra-scale bidirectional transformer does not inherently encode positional information, stacking multiple transformer layers can dilute the relative positional signals among tokens. To mitigate this issue, we first add positional encodings to the token embeddings after each transformer layer. In addition, we augment $\mathbf{M}_{k}$ with a position-aware soft masking matrix $\mathbf{M}^{p}_{k}\in\mathbb{R}^{s_{k}\times s_{k}}$, which is derived from the coordinate-based absolute positional embedding matrix $\mathbf{P}_{k}\in\mathbb{R}^{s_{k}\times d}$:
$\mathbf{M}^{p}_{k}$ is a symmetric matrix, where each entry $\mathbf{M}^{p}_{k}[i,j]\in(0,1)$ encodes the soft relative position information between points $i$ and $j$. This raises a key question: how can we derive the positional embedding $\mathbf{P}_{k}$ for each scale when no explicit 3D geometry is available at this stage? Our solution is an intermediate-structure decoding strategy. Specifically, we apply the decoder $D(\cdot)$ to reconstruct the intermediate structure $\mathbf{X}_{k}$ using all ground-truth tokens up to step $k$:
$\mathbf{X}_{k}=D(\sum_{m=1}^{k}\phi_{m}(\operatorname{upsampling}(\mathbf{z}_{m},s_{m}))),$ (9)
where $(\mathbf{z}_{1},\dots,\mathbf{z}_{k})=\operatorname{lookup}(Z,(q_{1},\dots,q_{k}))$, $\mathbf{X}_{k}\in\mathbb{R}^{s_{k}\times 3}$ provides the coordinate information used to compute the positional embedding $\mathbf{P}_{k}$. $\mathbf{P}_{k}$ is derived as an absolute positional encoding based on the 3D coordinates $\mathbf{X}_{k}$ using trigonometric functions (e.g. $\sin$ and $\cos$). During inference, the predicted token $\hat{q}_{k}$ is used to obtain $\mathbf{\hat{z}}_{k}$ in Eq. 9, rather than the ground-truth token $q_{k}$, which is not available at test time. The detailed derivation of $\mathbf{P}$ is provided in the Appendix 7. Note that any positional encoding based on token indices should not be applied, as it would violate the permutation equivariance property.
The prediction of each token $\hat{q}_{k}^{i}$ is evaluated using the cross-entropy (CE) loss $\mathcal{L}^{i}_{k}=\operatorname{CE}(\hat{q}_{k}^{i},q^{i}_{k})$. We first compute intra-scale loss $\mathcal{L}_{k}=\frac{1}{s_{k}}\sum_{i=1}^{s_{k}}\mathcal{L}^{i}_{k}$ and then compute inter-scale loss $\mathcal{L}_{\text{total}}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{k}$. The second stage training architecture is illustrated in Figure 3 (b). We provide a theoretical analysis of the permutation invariance of PointNSP ’s distribution modeling in Appendix 6.
DPM [DPM] Diffusion 76.42 86.91 60.05 74.77 68.89 79.97 68.45 80.55
PVD [PVD] Diffusion 73.82 64.81 56.26 53.32 54.55 53.83 61.54 57.32
LION [LION] Diffusion 72.99 64.21 55.67 53.82 53.47 53.21 61.75 57.59